OpenStudy (poof):

Need help on learning how to do this problem. Solve the Inequality for x. 5 - 3/2x >= 1/3

OpenStudy (akashdeepdeb):

Is this what the inequality looks like? \[5-\frac{3}{2}x \geq \frac{1}{3}\]

OpenStudy (poof):

Yes

OpenStudy (akashdeepdeb):

Awesome. So, since this is an inequality, you've got to rearrange some terms to find the few (or a lot of) values of x that satisfy it. Let's say you have an inequality like: \[x + a \leq -b + c\] How do you solve something like this? Well, you follow similar rules you use to solve normal equations and find the solution (or the set of solutions here) for x. So, you basically rearrange the terms: \[x \leq c-b-a\] And so, we have for any value of a, b, c x is less than c - b - a. Does this make sense?

OpenStudy (poof):

What you're trying to say is that we need to re-arrange it by trying to isolate the X?

OpenStudy (akashdeepdeb):

Exactly! Try doing that and ask me for help if you don't get something.

OpenStudy (poof):

What i'm having trouble with is fractions. I don't know how to deal with them in equations or if i'm dividing -3/2 first on both sides before subtracting 5. Am i going left to right or right to left. If i were to attempt it, it would look something like this.